# When Your Boss Is an AI Bot: Exploring Opportunities and Risks of Manager Clone Agents in the Future Workplace

### 24 Sep 2025 v2

@@ 4.2. Design Fiction Workshops
Each design fiction workshop lasted approximately 120 minutes, following a structured sequence of activities designed to elicit both imaginative and critical reflections on Manager Clone Agents in future teams and organizations. To mitigate potential power asymmetries between managers and workers, we ensured that each participant had dedicated turns to speak during group activities, explicitly invited quieter participants to share their views. Structured facilitation further emphasized that both managerial and worker perspectives carried equal value, reducing the likelihood that managerial voices would dominate. IRB approval was obtained prior to the workshops; all participants provided informed consent, and each received \$30 gift card after participation.
@@
Building on these examples, participants individually brainstormed how their cases might unfold differently if a manager had been replaced by a Manager Clone Agent. Managers were asked to adopt the perspective of a leader delegating authority to a Manager Clone Agent, while workers were asked to imagine themselves being directed or supported by such an agent. This activity helped situate speculative thinking in the specific dynamics of authority, oversight, and collaboration familiar to each role.
@@
Each participant then selected one scenario to develop in greater depth. They spent 15 minutes writing a short design fiction narrative describing how future workplace practices, relationships, and outcomes would unfold when Manager Clone Agents were involved. Participants were encouraged to build directly on the examples they had shared earlier, but were also free to expand creatively and develop entirely new directions. This balance ensured that narratives remained grounded in lived experience while allowing space for imaginative exploration of near-future possibilities. It is noted that these design fictions are based on the imaginaries from our participants, instead of existing Manager Clone Agents deployments.
@@
After writing, participants took turns presenting their stories to the group. Following each presentation, other participants engaged in a role-play discussion by responding as if they were members of that scenario, either managers or workers based on their own roles, reflecting on how they would experience and react to the presence of a Manager Clone Agent. This process encouraged participants to inhabit perspectives beyond their own and to surface tensions between managerial and worker vantage points.
@@ 5. Data Analysis
We analyzed the workshop materials using reflexive thematic analysis (Braun and Clarke, 2021, 2019). Our dataset included participants’ written design fiction stories, notes taken during scenario brainstorming, and transcripts from group discussions and round-table synthesis. All workshops were audio-recorded and transcribed verbatim. To preserve the richness of participants’ accounts, we treated both fictional narratives and reflective commentary as complementary forms of data: the stories captured imaginative possibilities, while the discussions revealed participants’ evaluations, concerns, and reasoning.
@@
The analysis followed an iterative, multi-stage process. First, two researchers independently read through all transcripts and narratives to gain familiarity with the data and to note initial impressions. Second, we conducted open coding at the sentence and paragraph level, tagging both descriptive content (e.g., tasks automated, benefits cited, risks identified) and relational dynamics (e.g., authority, trust, recognition). Third, we collaboratively grouped codes into higher-level categories that captured recurring patterns across workshops. Throughout this process, we regularly revisited raw data to ensure that emerging categories remained grounded in participants’ accounts.
In the final stage, we synthesized categories into three overarching themes that structured our findings: (1) the supportive roles participants envisioned for Manager Clone Agents (see section 6), (2) the risks and tensions they anticipated at individual, interpersonal, and organizational levels (see section 7), and (3) the future design directions they proposed (see section 8). These themes reflected both convergence and divergence in perspectives across managers and workers, highlighting how Manager Clone Agents were imagined not only as technical systems but as relational actors within organizational life. These findings themes are also summarized in Figure 2 for reference.
@@
To enhance trustworthiness, the research team engaged in regular debriefing sessions to challenge assumptions and refine interpretations. Disagreements were resolved through discussion, and multiple iterations of coding and theme development were used to balance consistency with openness to new insights. Following a traditional HCI qualitative analysis approach (e.g., (Xiao et al., 2025b; Fan et al., 2025)), we did not assess statistical reliability or validity; instead, our reflexive thematic analysis emphasized depth of interpretation, contextual grounding, and transparency in the analytic process.
@@
6. Findings: Empowering Teamwork: Four Potential Supportive Roles of the Manager Clone Agents (RQ1)
Our findings reveal that participants envisioned Manager Clone Agents to empower future teamwork across four distinct supportive roles. As a Proxy Presence, they were expected to mitigate the costs of managerial absence by sustaining responsiveness and symbolic availability when managers were busy and not available (see subsection 6.1). Acting as an Information Conveyor Belt, they were imagined to streamline communication flows, reduce distortion, and surface grassroots concerns that might otherwise be lost in hierarchies (see subsection 6.2). As a Productivity Engine, they were expected to expand managers’ capacity by automating routine tasks while also freeing cognitive bandwidth for more strategic work (see subsection 6.3). Finally, as an Amplifier of Leadership, they were envisioned to extend the manager’s coaching, feedback, and mentoring to a broader set of workers, compensating for the limits of managerial time and attention and magnifying guidance that would otherwise remain scarce (see subsection 6.4).
6.1. A Proxy Presence: Always Present, Always on Call
@@
Many managers and worker participants in our study emphasized that the burdensome nature of leadership responsibilities often makes physical presence difficult. M6 described how managers like him are constantly on the move: “meeting clients, traveling across regions, and balancing competing responsibilities.” He added that this constant mobility left him mentally and physically exhausted, as he was expected to maintain visibility and responsiveness while being pulled across multiple sites of work. Other managers echoed this sentiment, noting that such demands delayed responses and reduced availability. Several workers, including W4, W6, and W8, also described feeling hesitant to approach managers who were perpetually “elsewhere”, worrying that asking for help would add to their burden.
@@
M9, a marketing manager with 5–10 years of experience, illustrated this possibility in the speculative design story she wrote. In her scenario, a family emergency coincided with an important cross-timezone stakeholder meeting. Rather than canceling, she imagined sending her avatar, having pre-uploaded key talking points so that “the discussion could still move forward even if I wasn’t there in person.”
@@
Beyond crisis situations, participants also saw Manager Clone Agents as tools for more flexible participation. By mediating communication asynchronously, they allowed both managers and workers to engage at their own pace rather than in real time. W8 explained: “With the Manager Clone Agent standing in, it’s the same as leaving a voice message. I can just listen whenever I have time, until it’s the best time for me.”
@@
Additionally, compared with interacting directly with the real manager, some worker participants described feeling less pressure to give undivided attention when meeting with clone agents. As W8 put it, “If they’re talking about results, I’ll just leave it on in the background, like a Zoom meeting with camera and mic off.” This comment highlighted how the mediated presence of a Manager Clone Agent shifted expectations: workers could still access managerial input without the same demand for immediate, face-to-face attentiveness.
@@
Participants also noted that the symbolic presence of a Manager Clone Agent carries different weight than sending an assistant or proxy human. As M12 put it, “Even though it’s a clone, it’s your leader’s clone, not another person who might distort information.” Here, the symbolic value lies in the continuity of identity: the clone embodies the manager’s voice and authority, even if mediated, whereas an assistant is seen as a separate actor. M10 believed that, for broadcast-style meetings or official statements, this symbolic link made the clone agent feel more authoritative and representative than delegating to other staff. At the same time, participants emphasized that symbolic presence has limits. As M4 noted, the manager’s personal appearance still matters in high-stakes contexts where trust cannot be fully substituted by a clone.
6.2. An Information Conveyor Belt: Filtering Noise and Freeing Voices
@@
Participants repeatedly highlighted the communicative role of managers as conduits of information—“passing directives downward and carrying feedback upward, particularly for mid-level managers” as M6 described. They envisioned Manager Clone Agents as impartial conveyors that could transmit policies and feedback more clearly, reduce cognitive misalignment, and create safer channels for workers’ voices to be heard. Beyond individual exchanges, participants also imagined these agents improving large-scale information gathering, amplifying grassroots concerns that might otherwise be lost or filtered.
@@
In discussing impartial transmission, M10 noted that messages relayed through human intermediaries were “less likely to be impartial”, often shaped by bias, incomplete understanding, or social discomfort. Building on this, workers like W3 and W4, along with managers such as M12, imagined Manager Clone Agents as more neutral carriers. In their stories, agents could smooth over gaps, keep meaning intact, and reduce the confusion that often arises when policies and feedback travel through layers of hierarchy.
@@
Participants also reflected on how Manager Clone Agents might reduce the friction caused by cognitive misalignment between managers and workers. W4 recounted the difficulty of explaining her visa situation to a manager unfamiliar with international work rules: “Earlier I discussed with my manager how I could extend my OPT. Since he is not an international worker, I needed to provide him with some context and continue to tell him what my timeline looks like, and he has to continue to report to his manager and human resources.” In her view, the problem was not unwillingness but the sheer effort required to keep restating context across different levels of hierarchy. A Manager Clone Agent, participants suggested, could retain and transmit such contextual knowledge on behalf of the worker, ensuring that details are not lost each time the message is passed upward, while also sparing workers from the exhausting cycle of repeated clarification.
@@
Another recurring theme was the potential to free voices that might otherwise be silenced by hierarchy. M10 observed that “in the workplace, workers may have some requests and not dare say them directly.” W9 in the same workshop echoed this, imagining Clone Agents as mediators that could create safer channels for raising concerns. In one speculative story, M2 described recruitment scenarios where applicants could ask sensitive questions more comfortably to an agent than to a manager, since the agent could respond with patience and without judgment, and applicants would not feel embarrassed or disadvantaged for raising such issues.
@@
Finally, participants described how Manager Clone Agents might improve large-scale information collection. M12 imagined an agent preparing for a general assembly by first meeting with smaller groups to surface pressing issues: “To prepare for a general assembly, an Agent could meet with smaller groups (departments, floors, etc.) in order to figure out what pressing issues should be addressed.” W10, in the same workshop, agreed that this would make workers feel heard, contrasting it with surveys: “If you have the Manager Clone Agent, you will feel like someone is listening to you, and they will do something about it.”
6.3. A Productivity Engine: Automating Tasks and Augmenting Intelligence
Participants imagined Manager Clone Agents as productivity engines that could expand managerial capacity in two complementary ways: by automating repetitive, rule-based tasks and by extending cognitive functions that support decision-making. These dual capacities suggested a vision of leadership freed from low-level burdens and better equipped to focus on strategy, growth, and human relationships.
@@ 6.3. A Productivity Engine: Automating Tasks and Augmenting Intelligence
Many participants described how Manager Clone Agents could absorb what M7 called “high-frequency, low-risk, and rule-based” processes that otherwise drained managerial time and attention. Routine approvals, such as “approving invoices” (M4), “verifying expense claims” (M7), “processing leave” (W1, M6), or “signing off on weekly timesheets” (W6), were repeatedly mentioned as tasks that could be offloaded.
As M7 explained, these workflows were “essentially repetitive approvals that follow preset rules; having an agent check and approve them would save enormous effort without shifting responsibility.” Others emphasized that such automation could bring flexibility, with W1 noting that conflicts or sick leave could be handled more smoothly if agents were able to reschedule meetings or process minor requests.
In more complex contexts, participants even imagined agents carrying through contract reviews or resource allocations once decisions had already been settled. M7, for instance, told a story where “Party B mailed back the stamped contract (which had already been revised several times). They hope we approve and stamp it promptly and release the funds. At this point, the AI agent, based on the previously discussed details […] then proceeds with the stamping process, seeks finance approval, and completes the disbursement procedures.” She estimated that such automations could offload up to “10–15% of the workload.” Some, like M10, went further, imagining multi-agent setups where a Manager Clone Agent orchestrated specialized sub-agents for scheduling, finance, or reporting, thereby scaling efficiency beyond what one manager could achieve.
Participants further emphasized that agents might enhance fairness and accuracy in managerial decisions. M6 believed that by maintaining records and reducing reliance on subjective memory, agents could bring greater impartiality to areas such as task allocation or performance evaluation. M7 imagined them “reducing errors caused by fatigue, standardizing compliance checks,” while M3 suggested they could support “fairer, data-driven judgments.” At the same time, there are some cautions. M1 and W3 worried that delegating too much discretion could backfire if managers later overturned an agent’s decision, eroding credibility. For most of the participants, the promise of cognitive support had to be carefully balanced against the risk of eroding managerial authority.
6.4. An Amplifier of Leadership: Scaling Support, Enabling Growth

### 29 Jan 2026 v3

@@ 4.2. Design Fiction Workshops
Each design fiction workshop lasted approximately 120 minutes, following a structured sequence of activities designed to elicit both imaginative and critical reflections on Manager Clone Agents in future teams and organizations. Aligned with design fiction and speculative design traditions in HCI (Dunne and Raby, 2024; Wong and Khovanskaya, 2018), we convened managers and workers in a single workshop and created room for debate about socio-technical values and power dynamics. To offset the power asymmetries between managers and workers, we ensured that each participant had dedicated turns to speak during group activities, and rotated the speaking order of managers and workers in each speaking round. In addition, to reducing the likelihood that managerial voices would dominate, we explicitly encouraged workers in weaker positions of power to comment on or challenge the managers’ concepting. IRB approval was obtained prior to the workshops; all participants provided informed consent, and each received \$30 gift card after participation.
@@
Building on these examples, participants individually brainstormed how their cases might unfold differently if a manager had been replaced by a Manager Clone Agent. Managers were asked to adopt the perspective of a leader delegating authority to a Manager Clone Agent, while workers were asked to imagine themselves being directed by such an agent. This activity served as a recall scaffold that helped participants surface tacit power dynamics and conflicts to later reconfigure in fiction.
@@
Each participant then selected one scenario to develop in greater depth, but was also free to expand creatively and develop entirely new directions. They spent 15 minutes writing a short design fiction narrative describing how future workplace practices and relationships would unfold when Manager Clone Agents were involved. We prompted them to stretch or exceed present-day technological assumptions whenever their stories required it. Participants were encouraged to build directly on the examples they had shared earlier, but were also free to expand creatively and develop entirely new directions, including resistant or structurally disruptive possibilities. This balance ensured that narratives remained grounded in lived experience, while allowing space for imaginative exploration of alternative future possibilities. It is noted that these design fictions are based on the imaginaries of our participants, instead of existing Manager Clone Agents deployments.
@@
After writing, the participants took turns presenting their stories to the group. Following each presentation, other participants engaged in a role-play discussion by responding as if they were members of that scenario, either managers or workers based on their own roles, reflecting on how they would experience and react to the presence of a Manager Clone Agent. This process encouraged participants to inhabit perspectives beyond their own and to surface tensions between managerial and worker vantage points.
@@ 4.3. Data Analysis
We analyze the workshop materials using reflexive thematic analysis (Braun and Clarke, 2021, 2019). The dataset included participants’ written design fiction stories, notes from scenario brainstorming, and transcripts from group discussions and round-table synthesis. All workshops were audio-recorded and transcribed verbatim. To preserve the richness of participants’ accounts, we treated fictional narratives and reflective commentary as complementary forms of data: stories articulated imaginative possibilities, while discussions revealed participants’ interpretations, concerns, and reasoning about those imagined futures.
Our analytical approach was primarily inductive. Initial codes were generated bottom-up from participants’ narratives and discussions rather than derived from predefined theoretical categories, while still being informed by sensitizing concepts from CSCW/HCI research. This orientation allowed us to remain open to unexpected insights while ensuring that interpretations were grounded in participants’ meaning-making.
@@
The analysis proceeded iteratively. First, two researchers independently read through all transcripts and narratives to gain familiarity with the data and note initial impressions. Second, we conducted open coding at the sentence and paragraph level, tagging both descriptive content (e.g., tasks automated, benefits cited, risks identified) and relational dynamics (e.g., authority, trust, recognition). Third, we collaboratively grouped codes into higher-level categories that captured recurring patterns across workshops. Throughout this process, we repeatedly returned to the raw data to ensure that emerging categories remained grounded in participants’ accounts.
The final stage involved synthesizing categories into three overarching themes that structured our findings (see Figure 2) : (1) the supportive roles participants envisioned for Manager Clone Agents (see section 5), (2) the risks and tensions anticipated at individual, interpersonal, and organizational levels (see section 6), and (3) the future design directions participants proposed (see section 7). Although the themes align with the study’s research questions, this alignment reflects the shared grounding of both the questions and the themes in participants’ accounts. The research questions framed the scope of inquiry, but the themes themselves emerged inductively through the bottom-up clustering of categories rather than from a predefined coding frame.
Design fiction played a central role in shaping our analytic process. Because participants’ narratives intentionally exaggerated, extended, or reconfigured familiar workplace dynamics, the fictional mode acted as a probe that surfaced latent assumptions about authority, collaboration, delegation, and managerial presence. During coding, we treated fictional exaggerations not as predictions but as windows into participants’ organizational imaginaries, enabling us to identify concerns and relational patterns that may remain unarticulated in conventional interview settings.
@@
To enhance trustworthiness, the research team held regular debriefing sessions to challenge assumptions and refine interpretations. Disagreements were resolved through discussion and multiple iterations of coding and theme development were used to balance consistency with openness to new insights. Consistent with qualitative HCI practices (e.g., (Xiao et al., 2025; Fan et al., 2025)), we did not assess statistical reliability; instead, we emphasized interpretive depth, contextual grounding, and transparency in the analytic process.
@@
Conceptual framework of findings organized around three research questions. For RQ1 (opportunities), managers and workers envision four supportive roles for Manager Clone Agents: proxy presence (always present/on call), information conveyor (filtering noise and freeing voices), productivity engine (automating tasks and augmenting intelligence), and leadership amplifier (scaling support and enabling growth). For RQ2 (challenges/risks), anticipated tensions cluster at individual (accountability, career growth, irreplaceability), interpersonal (trust, authenticity, affective communication), and organizational levels (efficiency-first logics, flat hierarchies, weak community culture), shown as overlapping layers. For RQ3 (design considerations), a tiered pyramid highlights worker-oriented principles (fairness/equity/respect, tangible value, mutual consent, avoiding excessive control/surveillance), a call to revisit “enhance vs. replace” (extend capacity without replacing authority, retain human autonomy, support relationships), and tiered autonomy (representation → proactivity → delegation, with gradual trust-building and flexible configuration). The base emphasizes disruptive and unsettling alternative futures, posing critical “what if” scenarios (workers refusing clones, managers fading, and clones becoming tools of resistance) that shape future workplace visions.
5. Findings (RQ1): Empowering Teamwork: Four Potential Supportive Roles of the Manager Clone Agents
Our findings reveal that participants envisioned Manager Clone Agents to empower future teamwork across four distinct supportive roles. As a Proxy Presence, they were expected to mitigate the costs of managerial absence by sustaining responsiveness and symbolic availability when managers were busy and not available (see subsection 5.1). Acting as an Information Conveyor, they were imagined to streamline communication flows, reduce distortion, and surface grassroots concerns that might otherwise be lost in hierarchies (see subsection 5.2). As a Productivity Engine, they were expected to expand managers’ capacity by automating routine tasks while also freeing cognitive bandwidth for more strategic work (see subsection 5.3). Finally, as an Amplifier of Leadership, they were envisioned to extend the manager’s coaching, feedback, and mentoring to a broader set of workers, compensating for the limits of managerial time and attention and magnifying guidance that would otherwise remain scarce (see subsection 5.4).
5.1. A Proxy Presence that Preserves Managers’ Continuity and Symbolic Authority
@@
Many managers and worker participants in our study emphasized that the burdensome nature of leadership responsibilities often makes physical presence difficult. M10 described how managers like him are constantly on the move: “meeting clients, traveling across regions, and balancing competing responsibilities.” He added that this constant mobility left him mentally and physically exhausted, as he was expected to maintain visibility and responsiveness while being pulled across multiple sites of work. Other managers echoed this sentiment, noting that such demands delayed responses and reduced availability. Several workers, including W8, W12, and W15, also described feeling hesitant to approach managers who were perpetually “elsewhere,” worrying that asking for help would add to their burden.
@@
M17, a marketing manager with 5–10 years of experience, illustrated this possibility in the speculative design story she wrote. In her scenario, a family emergency coincided with an important cross-timezone stakeholder meeting. Rather than canceling, she imagined sending her avatar, having pre-uploaded key talking points so that “the discussion could still move forward even if I wasn’t there in person.”
@@
Beyond crisis situations, participants also saw Manager Clone Agents as tools for more flexible participation. By asynchronously mediating communication, they allowed both managers and workers to engage at their own pace rather than in real time. W15 explained: “With the Manager Clone Agent standing in, it’s the same as leaving a voice message. I can just listen whenever I have time, until it’s the best time for me.”
@@
Additionally, compared with interacting directly with the real manager, some worker participants described feeling less pressure to give undivided attention when meeting with clone agents. As W15 put it, “If they’re talking about results, I’ll just leave it on in the background, like a Zoom meeting with camera and mic off.” This comment highlighted how the mediated presence of a Manager Clone Agent shifted expectations: workers could still access managerial input without the same demand for immediate, face-to-face attentiveness.
@@
Participants also noted that the symbolic presence of a Manager Clone Agent carries different weight than sending an assistant or proxy human. As M21 put it, “Even though it’s a clone, it’s your leader’s clone, not another person who might distort information.” Here, the symbolic value lies in the continuity of identity: the clone embodies the manager’s voice and authority, even if mediated, whereas an assistant is seen as a separate actor. M18 believed that, for broadcast-style meetings or official statements, this symbolic link made the clone agent feel more authoritative and representative than delegating to other staff. At the same time, participants emphasized that symbolic presence has limits. As M6 noted, the manager’s personal appearance still matters in high-stakes contexts where trust cannot be fully replaced by a clone.
5.2. An Information Conveyor that Clarifies Directives and Surfaces Grassroots Feedback
@@
Participants repeatedly highlighted the communicative role of managers as conduits of information—“passing directives downward and carrying feedback upward, particularly for mid-level managers,” as M10 described. They envisioned Manager Clone Agents as impartial conveyors that could transmit policies and feedback more clearly, reduce cognitive misalignment, and create safer channels for workers’ voices to be heard. Beyond individual exchanges, participants also imagined these agents improving large-scale information gathering, amplifying grassroots concerns that might otherwise be lost or filtered.
@@
In discussing impartial transmission, M18 noted that messages relayed through human intermediaries were “less likely to be impartial”, often shaped by bias, incomplete understanding, or social discomfort. Building on this, workers like W7 and W8, along with managers such as M21, imagined Manager Clone Agents as more neutral carriers. In their stories, agents could smooth over gaps, keep meaning intact, and reduce the confusion that often arises when policies and feedback travel through layers of hierarchy.
@@
Participants also reflected on how Manager Clone Agents might reduce friction caused by cognitive misalignment between managers and workers. W8 recounted the difficulty of explaining her visa situation to a manager unfamiliar with international work rules: “Earlier I discussed with my manager how I could extend my OPT. Since he is not an international worker, I needed to provide him with some context and continue to tell him what my timeline looks like, and he has to continue to report to his manager and human resources.” In her view, the problem was not unwillingness but the sheer effort required to keep restating context across different levels of hierarchy. A Manager Clone Agent, participants suggested, could retain and transmit such contextual knowledge on behalf of the worker, ensuring that details are not lost each time the message is passed upward, while also sparing workers from the exhausting cycle of repeated clarification.
@@
Another recurring theme was the potential to free voices that might otherwise be silenced by hierarchy. M18 observed that “in the workplace, workers may have some requests and not dare to say them directly.” W20 in the same workshop echoed this, imagining Clone Agents as mediators that could create safer channels for raising concerns. In one speculative story, M4 described recruitment scenarios where applicants could ask sensitive questions more comfortably to an agent than to a manager, since the agent could respond with patience and without judgment, and applicants would not feel embarrassed or disadvantaged for raising such issues.
@@
Finally, participants described how Manager Clone Agents might improve large-scale information collection. M21 imagined an agent preparing for a general assembly by first meeting with smaller groups to surface pressing issues: “To prepare for a general assembly, an Agent could meet with smaller groups (departments, floors, etc.) to figure out which pressing issues should be addressed.” W23, in the same workshop, agreed that this would make workers feel heard, contrasting it with surveys: “If you have the Manager Clone Agent, you will feel like someone is listening to you and they will do something about it.”
5.3. A Productivity Engine that Prevents Fatigue and Bias, Freeing Leaders for Strategic Work
Participants expected Manager Clone Agents to improve productivity in two directions. On the one hand, agents can take over repetitive, rule-based work and prevent mistakes that come from human limits such as personal preference, fatigue, or bias. On the other hand, agents can strengthen managers’ decision quality by remembering more context and giving more consistent recommendations. This combination would shift managers’ effort away from low-level tasks toward strategic and relational responsibilities.
@@ 5.3. A Productivity Engine that Prevents Fatigue and Bias, Freeing Leaders for Strategic Work
Many participants described how Manager Clone Agents could absorb what M13 called “high-frequency, low-risk, and rule-based” processes that otherwise drained managerial time and attention. Routine approvals, such as “approving invoices” (M6), “verifying expense claims” (M13), “processing leave” (W1, M10), or “signing off on weekly timesheets” (W12), were repeatedly mentioned as tasks that could be offloaded.
Agents were also expected to prevent performance drop-offs caused by human bias. M21 noted that “people may judge, mislead or distort information, whereas a system is imagined as less prone to such flaws”. M5 described that “leaders interviewing many candidates can lose focus and forget earlier details; an agent’s steadier attention would catch inconsistencies that humans miss when attention is not concentrated”.
Beyond avoiding failure, agents were expected to push performance beyond human capacity. Participants emphasized machine strengths like “deeper contextual memory” (M19) and instant access to “domain or internet knowledge” (W11). M19 suggested that “a CEO’s repeated all-hands or cultural onboarding could be delivered by an agent who might speak even better than the CEO, freeing leaders from repeating the same material” (M19). W11 similarly imagined agents searching for online information rapidly to improve proposal revisions when leaders are busy or outside their expertise (W11).
Finally, by compressing details into quick takeaways, agents could let leaders “review in five minutes and spend time on higher-level judgment and abstract decisions” (M9). But participants also warned that misalignment between agent actions and leader intent can create coordination costs or reversal risks. M2 feared that “even small deviations could lead to ‘180-degree’ execution swings and wasted meetings if the leader overturns the agent later” (M2). Others noted that “agents still miss human grayspace and tacit intentions in scheduling or tradeoffs” (W8).
5.4. An Leadership Amplifier that Scales Mentorship and Allows for Autonomy and Growth